OpenStudy (anonymous):

Factor the trinomial. 5t^2-23t-10

OpenStudy (anonymous):

(5t+2)(t-5)

OpenStudy (anonymous):

\[5\cdot (-10)=-50\]so i need to think of two numbers that multiply to -50, but add to -23. That would be -25 and 2. So now i have:\[5t^2-25t+2t-10=5t(t-5)+2(t-5)\]\[=(5t+2)(t-5)\]

OpenStudy (anonymous):

thanks guys